# Topological restrictions on relatively Anosov representations

### v1

@@ 1. Introduction
In this paper, we study topological restrictions on relatively hyperbolic groups admitting relatively Anosov representations into $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$. Relatively Anosov representations, introduced in [KL23] and [Zhu21] (see also [ZZ23] and [Wan23]), provide a higher-rank analogue of geometrically finite representations into rank-one Lie groups. The construction and theory of these representations is inspired by the theory of Anosov representations, which were introduced by Labourie [Lab06] and Guichard–Wienhard [GW12] and provide a higher-rank analogue of convex cocompact representations into rank-one Lie groups.
@@
In many cases, Anosov representations come from convex cocompact actions on properly convex domains in real projective spaces [DGK17, Zim21]. Similarly, given any relatively Anosov representation $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{PGL}}_{d}(\operatorname{\mathbb{R}})$ whose image preserves a properly convex domain $\Omega\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d})$, there is an action of $\rho(\Gamma)$ on a properly convex domain $\Omega^{\prime}\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ which is geometrically finite in an appropriate sense ([ZZ22, Prop. 1.13], see also [IZ23].)
@@
Many examples of Anosov representations come from embeddings of rank-one lattices—or more generally convex cocompact subgroups of rank-one Lie groups—into larger Lie groups and deformations of these embeddings. There are other sources of Anosov representations, for example Coxeter groups [DGKLM], fundamental groups of certain Gromov–Thurston manifolds [Kap07], and, more broadly, cubulated hyperbolic groups [DFWZ]. Nevertheless, most of these other domain groups are only known to admit Anosov representations into Lie groups of high dimension relative to the complexity of the domain group.
@@
Restrictions on Anosov representations were established in [Tso20, CT20]. In essence, these results state that restricting the complexity of the domain hyperbolic group $\Gamma$ relative to the complexity of the target Lie group $\mathsf{G}$ constrains what the domain group can be. Here, complexity of the domain can be measured in terms of cohomological dimension, and complexity of the target in terms of $\dim\mathsf{G}$ and/or the size of the parabolic subgroup $\mathsf{P}<\mathsf{G}$ with respect to which we have the Anosov condition. Very informally speaking, the smaller the Lie group and/or the larger the parabolic subgroup, the smaller the resulting flag space $\mathsf{G}/\mathsf{P}$, and the less complexity the limit maps associated to the Anosov representation can exhibit. When the target group is $\operatorname{\mathsf{SL}}_{3}(\operatorname{\mathbb{R}})$, for instance, the only groups admitting Anosov representations are surface groups and free groups. Here and below, by a surface group, we mean the fundamental group of a closed hyperbolic surface. When the target group is $\operatorname{\mathsf{SL}}_{4}(\operatorname{\mathbb{R}})$, any group admitting a projective Anosov representation is a convex cocompact subgroup of $\operatorname{\mathsf{PO}}(1,3)$.
@@
Analogously, many known examples of relatively Anosov representations originate from geometrically finite subgroups of rank-one Lie groups. There are other known examples, for instance free products of nilpotent factors [ZZ22, Prop. 1.19]. However, as in the Anosov case, all of the known examples where the domain groups are not geometrically finite subgroups of rank-one Lie groups require the target Lie groups to have high dimension.
@@
We can thus ask if the results in [Tso20, CT20] have relative analogues. In other words, what restrictions are placed on a relatively hyperbolic group if we ask for it to admit a relatively Anosov representation into a fixed semisimple Lie group, with respect to a fixed parabolic subgroup?
@@
We remark that the characterization of domain groups of $k$-Anosov representations into $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ for odd $1\leq k\leq m$ was obtained independently in [DGR23] and [PT23].
@@
Note that in the case where $\Gamma$ has cohomological dimension at least 3, then we will see, from Theorem 6.1, that $\Gamma$ must in fact be the fundamental group of a closed hyperbolizable 3-manifold. This is a particular instance of case (3) of Theorem 1.4. Without the bound on the cohomological dimension, in contrast to the non-relative case in [CT20, §4], we are not able to conclude that the 3-manifolds which appear in case (2) are necessarily hyperbolizable: there are many more 3-manifolds with fundamental groups which are hyperbolic relative to free abelian subgroups.
@@
Theorem 1.6 (Theorem 5.3).
Let $d,k\in\operatorname{\mathbb{N}}$ with $d\geq 2k+1$ and $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $(\Gamma,\operatorname{\mathcal{P}})$ be a relatively hyperbolic group whose Bowditch boundary $\partial(\Gamma,\operatorname{\mathcal{P}})$ contains properly a sphere of dimension $r_{\operatorname{\mathbb{K}}}(d-k)-1$ (where $r_{\operatorname{\mathbb{R}}}=1$ and $r_{\operatorname{\mathbb{C}}}=2$). Then there is no relatively $k$-Anosov representation $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{K}})$.
@@
Theorem 1.5 is analogous to Theorem 1.3 in Canary–Tsouvalas [CT20]. The result in Canary–Tsouvalas is formulated in terms of cohomological dimension, making use of the close relation between the cohomological dimension of a hyperbolic group $\Gamma$ and the topological dimension of its Gromov boundary $\partial_{\infty}\Gamma$ (due to Bestvina–Mess [BM91]): $\dim(\partial_{\infty}\Gamma)=\mathrm{cd}(\Gamma)-1$.
In the relative case, Manning–Wang have results partially generalizing this relation: if $\Gamma$ is a torsion-free group which is hyperbolic relative to nilpotent groups (so that $\Gamma$ is of type $F$, by work of Dahmani [Dah03]), and $\mathrm{cd}(\Gamma)<\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})$, then $\dim(\partial(\Gamma,\operatorname{\mathcal{P}}))=\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})-1=\mathrm{cd}(\Gamma)$ [MW20, Thm. 5.1]. Conjecturally, the first equality holds without the additional assumption that $\mathrm{cd}(\Gamma)<\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})$.
In terms of cohomological dimension, we obtain restrictions when $\Gamma$ has high cohomological dimension (relative to the dimension of the target) analogous to [CT20, Thm. 1.5]. A representation $\rho\colon\Gamma\rightarrow\operatorname{\mathsf{SL}}_{d}(\mathbb{R})$ is called a Benoist representation if $\rho$ has finite kernel and $\rho(\Gamma)$ preserves and acts properly discontinuously and cocompactly on a strictly convex domain of $\mathbb{P}(\mathbb{R}^{d})$.
@@
We recall that Benoist proved in [Ben04] that a discrete subgroup $\mathsf{\Gamma}$ of $\mathsf{PGL}_{d}(\mathbb{R})$, $d\geq 3$, which preserves and acts cocompactly on strictly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$, is necessarily Gromov hyperbolic and its boundary $\partial\Omega$ is of class $\mathcal{C}^{1}$. Moreover, by [GW12, Prop. 6.1], the inclusion representation $\Gamma\xhookrightarrow{}\mathsf{GL}_{d}(\mathbb{R})$ is $1$-Anosov.
@@ 2.1. Convergence groups and relatively hyperbolic groups
We say that a group $\Gamma$ acts on a compact Hausdorff metrizable space $M$ as a convergence group if for every sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of distinct elements in $\Gamma$ there exists a subsequence $(\gamma_{k_{n}})_{n\in\mathbb{N}}$ and points $x,y\in M$ such that $(\gamma_{k_{n}}|_{M\smallsetminus\{y\}})_{n\in\mathbb{N}}$ converges locally uniformly to the constant map $x$. For a general reference on convergence groups, see e.g. [Tuk94].
Given $\Gamma$ acting on $M$ as a convergence group, an element $\gamma\in\Gamma$ is called loxodromic if it has infinite order and fixes exactly two points in $M$. The limit set $\Lambda_{\Gamma}\subset M$ is the set of points $x\in M$ where there exist $y\in M$ and a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ in $\Gamma$ where $(\gamma_{n}|_{M\smallsetminus\{y\}})_{n\in\mathbb{N}}$ converges locally uniformly to the constant map $x$. A subgroup $\Gamma_{0}$ of $\Gamma$ is called elementary it there exists a finite subset of $M$ with at most $2$ points fixed by $\Gamma_{0}$. If $\Gamma_{0}$ is non-elementary then $\Lambda_{\Gamma_{0}}$ is infinite, perfect and contains loxodromic elements, see [Tuk94, Thm. 2S + 2T].
It is a result of Bowditch that a relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ acts as a geometrically finite convergence group on its boundary $\partial(\Gamma,\operatorname{\mathcal{P}})$, with maximal parabolic subgroups given by the conjugates of subgroups in $\operatorname{\mathcal{P}}$ [Bow12, §6]. Indeed, there is a characterization of relatively hyperbolic groups in terms of geometrically finite convergence group actions [Yam04]. In particular, when $\Gamma$ is non-elementary, $\partial(\Gamma,\operatorname{\mathcal{P}})$ is uncountable and perfect. For a general reference on relatively hyperbolic groups, see e.g. [Bow12] or [Wei22].
@@ 2.2. Convergence groups acting on the circle
Theorem 2.1 (Gabai [Gab92] +Tukia [Tuk88]).
@@
We can combine this with Dunwoody’s accessibility theorem to obtain a statement about finite extensions of Fuchsian groups, whose proof we include here for completeness. For the notion of the fundamental group of a finite graph of groups we refer the reader to [Ser03].
@@
Let Γsuperscriptnormal-$\Gamma^{\prime}$ be a finitely generated discrete subgroup of $\operatorname{\mathsf{PSL}}_{2}(\operatorname{\mathbb{R}})$. Suppose that $1\to F\to\Gamma\to\Gamma^{\prime}\to 1$ is a short exact sequence of groups where $F$ is finite. Then $\Gamma$ contains a finite-index subgroup isomorphic to a free group or a surface group.
@@
Since $\Gamma^{\prime}$ acts properly discontinuously on the hyperbolic plane $\operatorname{\mathbb{H}}^{2}$, it is virtually a free or a surface group. If $\Gamma^{\prime}$ is virtually a surface group then $\Gamma$ is a hyperbolic group with $\partial_{\infty}\Gamma=\partial_{\infty}\Gamma^{\prime}\cong S^{1}$. By the work of Gabai [Gab92] it follows that $\Gamma^{\prime}$ is virtually a surface group.
Now suppose that $\Gamma^{\prime}$ is virtually free. Note that $\Gamma$ is a finitely presented group, so by Dunwoody’s accessibility theorem [Dun85], $\Gamma$ splits as the fundamental group of a finite graph of groups $\mathcal{G}$ with edge groups of at most one end. Note that since $\partial_{\infty}\Gamma^{\prime}\cong\partial_{\infty}\Gamma$ is totally disconnected, all vertex groups of $\mathcal{G}$ are virtually cyclic and thus $\Gamma$ itself is virtually free. ∎
@@ 2.3. Remarks on torsion
The proposition follows immediately from Gromov’s polynomial growth theorem [Gro81], since finite extensions of virtually nilpotent groups still have polynomial growth.
@@ 2.4. Splittings and boundaries of relatively hyperbolic groups
Theorem 2.7 ([Bow12, Prop. 10.1-3], see also [DG08, §3]).
@@
Suppose we have a non-trivial splitting of $\Gamma$ over finite groups relative to $\operatorname{\mathcal{P}}$. Then, from the discussion following Proposition 10.3 in [Bow12], there is a natural continuous inclusion of the boundary of each of the vertex subgroups (i.e. the limit set of each of these subgroups in $X(\Gamma,\operatorname{\mathcal{P}})\cup\partial(\Gamma,\operatorname{\mathcal{P}})$) into $\partial(\Gamma,\operatorname{\mathcal{P}})$. Indeed, each component of $\partial(\Gamma,\operatorname{\mathcal{P}})$ is either a single point or the boundary of an infinite non-peripheral vertex group. Singleton components are either fixed points of vertex groups which are peripheral, or else corresponds to ideal points of the tree corresponding to the splitting.
We remark that this is not the JSJ splitting over elementary subgroups given by Guirardel–Levitt [GL17] and described in Haulmark–Hruska [HH23]: for instance, the edge groups in the JSJ splitting are not in general finite. The JSJ splitting can be viewed as a further canonical maximal splitting of each of the vertex groups with connected boundary.

### v2

# rel_nonAnosov.bib

@@ 1. Introduction
In this paper, we study topological restrictions on relatively hyperbolic groups admitting relatively Anosov representations into $\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{R}})$. Relatively Anosov representations, introduced in [KapovichLeeb] and [reldomreps] (see also [ZZ1] and [Tianqi]), provide a higher-rank analogue of geometrically finite representations into rank-one Lie groups. The construction and theory of these representations is inspired by the theory of Anosov representations, which were introduced by Labourie [Lab] and Guichard–Wienhard [GW] and provide a higher-rank analogue of convex cocompact representations into rank-one Lie groups.
@@
In many cases, Anosov representations come from convex cocompact actions on properly convex domains in real projective spaces [DGK, Zimmer]. Similarly, given any relatively Anosov representation $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{PGL}}_{d}(\operatorname{\mathbb{R}})$ whose image preserves a properly convex domain $\Omega\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d})$, there is an action of $\rho(\Gamma)$ on a properly convex domain $\Omega^{\prime}\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ which is geometrically finite in an appropriate sense ([ZZ2, Prop. 1.13], see also [rAGF_Mitul].)
@@
Many examples of Anosov representations come from embeddings of rank-one lattices—or more generally convex cocompact subgroups of rank-one Lie groups—into larger Lie groups and deformations of these embeddings. There are other sources of Anosov representations, for example Coxeter groups [DGKLM], fundamental groups of certain Gromov–Thurston manifolds [KapovichGT], and, more broadly, cubulated hyperbolic groups [DFWZ]. Nevertheless, most of these other domain groups are only known to admit Anosov representations into Lie groups of high dimension relative to the complexity of the domain group.
@@
Restrictions on Anosov representations were established in [Tsouvalas, CanaryTsouvalas]. In essence, these results state that restricting the complexity of the domain hyperbolic group $\Gamma$ relative to the complexity of the target Lie group $\mathsf{G}$ constrains what the domain group can be. Here, complexity of the domain can be measured in terms of cohomological dimension, and complexity of the target in terms of $\dim\mathsf{G}$ and/or the size of the parabolic subgroup $\mathsf{P}<\mathsf{G}$ with respect to which we have the Anosov condition. Very informally speaking, the smaller the Lie group and/or the larger the parabolic subgroup, the smaller the resulting flag space $\mathsf{G}/\mathsf{P}$, and the less complexity the limit maps associated to the Anosov representation can exhibit. When the target group is $\operatorname{\mathsf{SL}}_{3}(\operatorname{\mathbb{R}})$, for instance, the only groups admitting Anosov representations are surface groups and free groups. Here and below, by a surface group, we mean the fundamental group of a closed hyperbolic surface. When the target group is $\operatorname{\mathsf{SL}}_{4}(\operatorname{\mathbb{R}})$, any group admitting a projective Anosov representation is a convex cocompact subgroup of $\operatorname{\mathsf{PO}}(1,3)$.
@@
Analogously, many known examples of relatively Anosov representations originate from geometrically finite subgroups of rank-one Lie groups. There are other known examples, for instance free products of nilpotent factors [ZZ2, Prop. 1.19]. However, as in the Anosov case, all of the known examples where the domain groups are not geometrically finite subgroups of rank-one Lie groups require the target Lie groups to have high dimension.
@@
We can thus ask if the results in [Tsouvalas, CanaryTsouvalas] have relative analogues. In other words, what restrictions are placed on a relatively hyperbolic group if we ask for it to admit a relatively Anosov representation into a fixed semisimple Lie group, with respect to a fixed parabolic subgroup?
@@
We remark that the characterization of domain groups of $k$-Anosov representations into $\operatorname{\mathsf{Sp}}_{2m}(\mathbb{R})$ for odd $1\leq k\leq m$ was obtained independently in [DGR], using results from [Dey-Borel], and in [PT].
@@
Note that in the case where $\Gamma$ has cohomological dimension at least 3, then we will see, from Theorem 6.1, that $\Gamma$ must in fact be the fundamental group of a closed hyperbolizable 3-manifold. This is a particular instance of case (3) of Theorem 1.4. Without the bound on the cohomological dimension, in contrast to the non-relative case in [CanaryTsouvalas, §4], we are not able to conclude that the 3-manifolds which appear in case (2) are necessarily hyperbolizable: there are many more 3-manifolds with fundamental groups which are hyperbolic relative to free abelian subgroups.
@@
Theorem 1.6.
Let $d,k\in\operatorname{\mathbb{N}}$ with $d\geq 2k+1$ and $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $(\Gamma,\operatorname{\mathcal{P}})$ be a relatively hyperbolic group whose Bowditch boundary $\partial(\Gamma,\operatorname{\mathcal{P}})$ properly contains a sphere of dimension $r_{\operatorname{\mathbb{K}}}(d-k)-1$ (where $r_{\operatorname{\mathbb{R}}}=1$ and $r_{\operatorname{\mathbb{C}}}=2$). Then there is no relatively $k$-Anosov representation $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\operatorname{\mathsf{SL}}_{d}(\operatorname{\mathbb{K}})$.
@@
Theorem 1.5 is analogous to Theorem 1.3 in Canary–Tsouvalas [CanaryTsouvalas]. The result in Canary–Tsouvalas is formulated in terms of cohomological dimension, making use of the close relation between the cohomological dimension of a hyperbolic group $\Gamma$ and the topological dimension of its Gromov boundary $\partial_{\infty}\Gamma$ (due to Bestvina–Mess [Bestvina-Mess]): $\dim(\partial_{\infty}\Gamma)=\mathrm{cd}(\Gamma)-1$.
In the relative case, Manning–Wang have results partially generalizing this relation: if $\Gamma$ is a torsion-free group which is hyperbolic relative to nilpotent groups (so that $\Gamma$ is of type $F$, by work of Dahmani [Dahmani_finite]), and $\mathrm{cd}(\Gamma)<\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})$, then $\dim(\partial(\Gamma,\operatorname{\mathcal{P}}))=\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})-1=\mathrm{cd}(\Gamma)$ [ManningWang, Thm. 5.1]. It is unclear whether the first equality holds without the additional assumption that $\mathrm{cd}(\Gamma)<\mathrm{cd}(\Gamma,\operatorname{\mathcal{P}})$, given recent results relating the cohomological dimension to a different compactification of a relatively hyperbolic group [Fukaya].
In terms of cohomological dimension, we obtain restrictions when $\Gamma$ has high cohomological dimension (relative to the dimension of the target) analogous to [CanaryTsouvalas, Thm. 1.5]. A representation $\rho\colon\Gamma\rightarrow\operatorname{\mathsf{SL}}_{d}(\mathbb{R})$ is called a Benoist representation if $\rho$ has finite kernel and $\rho(\Gamma)$ preserves and acts properly discontinuously and cocompactly on a strictly convex domain of $\mathbb{P}(\mathbb{R}^{d})$.
@@
We recall that Benoist proved in [BenoistI] that a discrete subgroup $\mathsf{\Gamma}$ of $\mathsf{PGL}_{d}(\mathbb{R})$, $d\geq 3$, which preserves and acts cocompactly on strictly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$, is necessarily Gromov hyperbolic and its boundary $\partial\Omega$ is of class $\mathcal{C}^{1}$. Moreover, by [GW, Prop. 6.1], the inclusion representation $\Gamma\xhookrightarrow{}\mathsf{GL}_{d}(\mathbb{R})$ is $1$-Anosov.
@@ 2.1. Convergence groups and relatively hyperbolic groups
We say that a group $\Gamma$ acts on a compact Hausdorff metrizable space $M$ as a convergence group if for every sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of distinct elements in $\Gamma$ there exists a subsequence $(\gamma_{k_{n}})_{n\in\mathbb{N}}$ and points $x,y\in M$ such that $(\gamma_{k_{n}}|_{M\smallsetminus\{y\}})_{n\in\mathbb{N}}$ converges locally uniformly to the constant map $x$. For a general reference on convergence groups, see e.g. [Tukia].
Given $\Gamma$ acting on $M$ as a convergence group, an element $\gamma\in\Gamma$ is called loxodromic if it has infinite order and fixes exactly two points in $M$. The limit set $\Lambda_{\Gamma}\subset M$ is the set of points $x\in M$ where there exist $y\in M$ and a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ in $\Gamma$ where $(\gamma_{n}|_{M\smallsetminus\{y\}})_{n\in\mathbb{N}}$ converges locally uniformly to the constant map $x$. A subgroup $\Gamma_{0}$ of $\Gamma$ is called elementary it there exists a finite subset of $M$ with at most $2$ points fixed by $\Gamma_{0}$. If $\Gamma_{0}$ is non-elementary then $\Lambda_{\Gamma_{0}}$ is infinite, perfect and contains loxodromic elements, see [Tukia, Thm. 2S + 2T].
It is a result of Bowditch that a relatively hyperbolic group $(\Gamma,\operatorname{\mathcal{P}})$ acts as a geometrically finite convergence group on its boundary $\partial(\Gamma,\operatorname{\mathcal{P}})$, with maximal parabolic subgroups given by the conjugates of subgroups in $\operatorname{\mathcal{P}}$ [Bowditch_relhyp, §6]. Indeed, there is a characterization of relatively hyperbolic groups in terms of geometrically finite convergence group actions [Yaman]. In particular, when $\Gamma$ is non-elementary, $\partial(\Gamma,\operatorname{\mathcal{P}})$ is uncountable and perfect. For a general reference on relatively hyperbolic groups, see e.g. [Bowditch_relhyp] or [Teddy_EGF].
@@ 2.2. Convergence groups acting on the circle
Theorem 2.1 (Gabai [Gabai] +Tukia [Tukia-Fuchsian]).
@@
We can combine this with Dunwoody’s accessibility theorem to obtain a statement about finite extensions of Fuchsian groups, whose proof we include here for completeness. For the notion of the fundamental group of a finite graph of groups we refer the reader to [Serre-trees].
@@
Let $\Gamma^{\prime}$ be a finitely generated discrete subgroup of $\operatorname{\mathsf{PSL}}_{2}(\operatorname{\mathbb{R}})$. Suppose that $1\to F\to\Gamma\to\Gamma^{\prime}\to 1$ is a short exact sequence of groups where $F$ is finite. Then $\Gamma$ contains a finite-index subgroup isomorphic to a free group or a surface group.
@@
Since $\Gamma^{\prime}$ acts properly discontinuously on the hyperbolic plane $\operatorname{\mathbb{H}}^{2}$, it is virtually a free or a surface group. If $\Gamma^{\prime}$ is virtually a surface group then $\Gamma$ is a hyperbolic group with $\partial_{\infty}\Gamma=\partial_{\infty}\Gamma^{\prime}\cong S^{1}$. By the work of Gabai [Gabai], it follows that $\Gamma$ is virtually a surface group.
Now suppose that $\Gamma^{\prime}$ is virtually free. Note that $\Gamma$ is a finitely presented group, so by Dunwoody’s accessibility theorem [Dun], $\Gamma$ splits as the fundamental group of a finite graph of groups $\mathcal{G}$ with edge groups of at most one end. Note that since $\partial_{\infty}\Gamma^{\prime}\cong\partial_{\infty}\Gamma$ is totally disconnected, all vertex groups of $\mathcal{G}$ are virtually cyclic and thus $\Gamma$ itself is virtually free. ∎
@@ 2.3. Remarks on torsion
The proposition follows immediately from Gromov’s polynomial growth theorem [Gromov-nil], since finite extensions of virtually nilpotent groups still have polynomial growth.
@@ 2.4. Splittings and boundaries of relatively hyperbolic groups
Theorem 2.7 ([Bowditch_relhyp, Prop. 10.1-3], see also [DahmaniGroves, §3]).
@@
Suppose we have a non-trivial splitting of $\Gamma$ over finite groups relative to $\operatorname{\mathcal{P}}$. Then, from the discussion following Proposition 10.3 in [Bowditch_relhyp], there is a natural continuous inclusion of the boundary of each of the vertex subgroups (i.e. the limit set of each of these subgroups in $X(\Gamma,\operatorname{\mathcal{P}})\cup\partial(\Gamma,\operatorname{\mathcal{P}})$) into $\partial(\Gamma,\operatorname{\mathcal{P}})$. Indeed, each component of $\partial(\Gamma,\operatorname{\mathcal{P}})$ is either a single point or the boundary of an infinite non-peripheral vertex group. Singleton components are either fixed points of vertex groups which are peripheral, or else corresponds to ideal points of the tree corresponding to the splitting.
We remark that this is not the JSJ splitting over elementary subgroups given by Guirardel–Levitt [GuirardelLevittJSJ] and described in Haulmark–Hruska [HaulmarkHruskaJSJ]: for instance, the edge groups in the JSJ splitting are not in general finite. The JSJ splitting can be viewed as a further canonical maximal splitting of each of the vertex groups with connected boundary.